# Joint Velocity-Growth Flow Matching for Single-Cell Dynamics Modeling

### v1

@@ Abstract
Learning the underlying dynamics of single cells from snapshot data has gained increasing attention in scientific and machine learning research. The destructive measurement technique and cell proliferation/death result in unpaired and unbalanced data between snapshots, making the learning of the underlying dynamics challenging. In this paper, we propose joint Velocity-Growth Flow Matching (VGFM), a novel paradigm that jointly learns state transition and mass growth of single-cell populations via flow matching. VGFM builds an ideal single-cell dynamics containing velocity of state and growth of mass, driven by a presented two-period dynamic understanding of the static semi-relaxed optimal transport, a mathematical tool that seeks the coupling between unpaired and unbalanced data. To enable practical usage, we approximate the ideal dynamics using neural networks, forming our joint velocity and growth matching framework. A distribution fitting loss is also employed in VGFM to further improve the fitting performance for snapshot data. Extensive experimental results on both synthetic and real datasets demonstrate that VGFM can capture the underlying biological dynamics accounting for mass and state variations over time, outperforming existing approaches for single-cell dynamics modeling.
@@ 1 Introduction
Inferring the latent dynamics of complex systems from sparse and noisy data is a fundamental challenge in science and engineering. In many domains, e.g., stock markets [1], climate systems [2], and biological processes [3, 4, 5, 6, 7, 8, 9, 10], continuous trajectories are rarely fully observed by sensors. Instead, cross-sectional snapshot data collected at discrete time points are commonly provided. This challenge is especially important in single-cell RNA sequencing [11, 12, 13, 14, 15], where destructive sampling yields unpaired population-level snapshots across time without having the tracked individual cell fates. Furthermore, due to the mass changes during cellular development or response processes, observed data often exhibits mass unbalancedness across time, violating mass conservation. Consequently, reconstructing the time-evolving, unnormalized density function from limited samples has become an important research problem and attracted increasing attention [13, 14, 16].
@@
Deep learning-based models for dynamics inference have demonstrated great potential [17, 3, 18, 7, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. These models typically employ ordinary or stochastic differential equations (ODEs or SDEs) parameterized by neural networks to approximate the velocity field governing density evolution. One class of methods, based on simulation [29, 17, 3, 30, 23, 19, 18, 22, 25], generates synthetic trajectories by feeding initial data through neural networks and solving ODEs or SDEs numerically, and compares simulated results with observations for loss computation. However, these simulation-based methods rely heavily on numerical solvers during training, which significantly increases computational cost. In high-dimensional settings, the enlarged search space further exacerbates instability, hindering scalability and convergence. Another line of research focuses on simulation-free approaches [31, 32, 33, 34, 21, 35, 36, 37, 38, 39, 40, 28, 41], where the velocity field is trained efficiently by constructing conditional probability paths without trajectory simulation, yielding better efficiency and stability in training compared with simulation approaches.
@@
Nevertheless, the simulation-free approaches mentioned above rely solely on the velocity field $v$, ignoring the unbalancedness of observed data, which can result in incorrect reconstruction of underlying dynamics or unsatisfying generation performance [25, 35, 40]. Distributional imbalance is a common phenomenon in single-cell dynamics, where cellular proliferation and death occur, thus necessitating the introduction of a growth term $g$ to allow for mass increasing or decreasing. To address this, [42, 43] proposed to jointly learn velocity $v$ and growth $g$ by minimizing the Wasserstein-Fisher-Rao (WFR) metric [44, 45]. However, it mathematically enforces $v=\nabla g$, which lacks clear mechanistic justification in biological systems. Other WFR-inspired approaches [22, 25] design separate neural networks for $v$ and $g$ and prioritize fitting the distributions. However, these methods require heavy simulation during training, limiting their scalability to high dimensions.
@@ 2 Background and Related Works
Given two distributions $p_{0}$ and $p_{1}$ on domain $\Omega$ satisfying $\int_{\Omega}p_{0}(\mathbf{x})\,\mathrm{d}x=\int_{\Omega}p_{1}(\mathbf{x})\,\mathrm{d}x$, the optimal transport [46] aims to find a joint distribution of $p_{0}$ and $p_{1}$, named coupling, such that the transport cost is minimized, formulated as the following optimization problem:
@@
where $c$ is the cost function. For a map $T$, we define $T_{\#}p_{0}(\mathbf{x})=\int_{\mathbf{x}^{\prime}:T(\mathbf{x}^{\prime})=\mathbf{x}}p_{0}(\mathbf{x}^{\prime})\mathop{}\!\mathrm{d}\mathbf{x}^{\prime}$. Let ${\mbox{P}}^{i}(\mathbf{x}_{0},\mathbf{x}_{1})$ $=\mathbf{x}_{i}$ for $i=0,1$, then we have ${\rm P}_{\#}^{i}\pi(\mathbf{x}_{i})=\int_{\Omega}\pi(\mathbf{x}_{0},\mathbf{x}_{1})\,\mathrm{d}\mathbf{x}_{i}$ for $i=0,1$. When $c(\mathbf{x}_{0},\mathbf{x}_{1})=\|\mathbf{x}_{0}-\mathbf{x}_{1}\|^{2}$, by Brenier’s theorem [47], the optimal coupling can be expressed as $\pi^{*}=(\mathrm{Id},T^{*})_{\#}p_{0}$, where $T^{*}$ is called the Monge map.
@@
Flow matching [31, 32, 33] is a simulation-free approach where the velocity field is trained efficiently by constructing conditional probability paths without trajectory simulation. Building on this foundation, further improvements have been achieved by incorporating optimal transport guidance [34, 21], considering unbalancedness [35], accounting for manifold structures [36], and approximating Schrödinger bridge via flow and score matching [39], all of which has been applied to model single-cell dynamics [21, 39, 36, 35]. Different from these methods, our method explicitly models the simultaneous matching of both the velocity field and the growth function, driven from our developed ideal joint state transition and mass growth dynamics, which does not require the construction of conditional probability paths like the above approaches.
@@
Data imbalance is prevalent across a variety of domains, such as image synthesis and protein generation, motivating the development of flow matching models that account for unbalanced distributions [35, 40, 48]. However, these methods often do not model growth functions, which are essential in single-cell contexts. In the single-cell domain, several approaches have been proposed to model growth dynamics [42, 43, 5, 22, 49], yet they either fail to leverage the informative variations in cell abundance observable from snapshot data [42, 43], or rely heavily on computationally expensive simulations [5, 22, 49], limiting their scalability and efficiency. Differently, our method explicitly learns the growth function from observed snapshot data, yielding better performance for single cells than these methods as shown in experiments.
@@ 3 Method
Given the snapshot population of single cells, this paper aims to build the dynamic trajectory of single cells. Considering the unbalancedness between snapshot populations due to the undergo cell proliferation or death, we aim to build a model that can transform between unbalanced distributions by jointly learning a velocity field for controlling state transition and a growth function for controlling mass variation. Towards this goal, we propose joint Velocity-Growth Flow Matching (VGFM) based on semi-relaxed optimal transport to learn the dynamics of single cells. As illustrated in Fig. 2, we first present a decoupled understanding of state transition and mass growth for unbalanced dynamics based on semi-relaxed optimal transport (Figs. 2 (b), (d)), upon which we build a dynamic process between unbalanced distributions (Fig. 2 (c) ). Finally, we propose the velocity and growth flow matching using the built dynamic process to learn the velocity and growth with samples (Fig. 2 (a)). Next, we first describe the unbalanced dynamics of single-cell data, and then discuss the details of each component of our method.
@@ 3.3 Velocity and Growth Flow Matching
[…]
@@
that can generalize to new samples. We first estimate $\pi^{*}$ and $T^{*}$ using samples. Given samples $\{\mathbf{x}_{0}^{i}\}_{i=1}^{n}$ of $p_{0}$ and $\{\mathbf{x}_{0}^{i}\}_{j=1}^{m}$ of $p_{1}$, We seek the optimal transport plan $\pi^{0\to 1}\approx\pi^{*}(\mathbf{x}_{0},\mathbf{x}_{1})$ by solving the entropy-regularized semi-relaxed transport problem using Sinkhorn-based algorithm [50, 51] as
@@
where $[\cdot]_{i}$ is the $i$-th element. The last term in Eq. (12) is because $p_{0}(\mathbf{x}_{0}^{i})=1$, for $i=1,\cdots,n$. In experiments, $\mathcal{L}_{\rm VGFM}$ can be implemented using mini-batch samples. Specifically, we sample $(i,j)$ from $\pi^{0\to 1}$ and $t$ from $\mathcal{U}(0,1)$, then calculate the loss in square brackets. To improve robustness, we add a Gaussian noise to $\mathbf{x}_{t}$ in experiments.
@@ 3.4 Training Process
Besides the flow matching loss in Eq. (12), we also employ distribution fitting loss to improve performance further. Next, we introduce the distribution fitting loss and our training algorithm.
To achieve better performance, we incorporate the Wasserstein distance between generated samples and observed samples as part of the loss function. Specifically, let $\mathcal{X}_{t}=\{\mathbf{x}^{i}_{t}\}_{i=1}^{N_{t}}$ for $t=0,1,\cdots,T-1$ denote the observed samples at different time points, and denote $p(\mathcal{X}_{t})=\sum_{i=1}^{N_{t}}\delta_{\mathbf{x}^{i}_{t}}$. We define $\phi_{v_{\theta}}:\mathbb{R}^{d}\times\mathcal{T}\to\mathbb{R}^{d|\mathcal{T}|}$ as the trajectory mapping function parameterized by the neural network $v_{\theta}$, which takes a given starting point as the initial condition and outputs particle coordinates at time indices $\mathcal{T}$ according to ODE dynamics, where $\mathcal{T}$ denotes a set of time steps. Given an initial set $\mathcal{X}_{0}$, the model $\phi_{v_{\theta}}$ predicts particle positions at future time points in $\mathcal{T}$. Specifically, the predicted samples are computed by applying the neural ODE to $\mathcal{X}_{0}$ over the time indices $\{1,\cdots,T-1\}$, i.e., $\hat{\mathcal{X}}_{1},\cdots,\hat{\mathcal{X}}_{T-1}=\phi_{v_{\theta}}(\mathcal{X}_{0},\{1,\cdots,T-1\})$. Similarly, we define $\phi_{g_{\omega}}$ as the particle weight mapping function parameterized by the neural network $g_{\omega}$, which takes the initial weight of particle $i$ as input and outputs the corresponding weight values at $\mathcal{T}$ under ODE dynamics. For simplicity of notation, we use $\hat{w}(\hat{\mathbf{x}})\mbox{ for }\hat{\mathbf{x}}\in\hat{\mathcal{X}}_{t}$ to denote the weight generated by $\phi_{g_{\omega}}$ corresponding to the particles in set $\hat{\mathcal{X}}$ which is generated by $\phi_{v_{\theta}}$ The distribution fitting loss is defined as the Wasserstein distance as
$\mathcal{L}_{\mathrm{OT}}(\theta,\omega)=\sum_{t=1}^{T-1}\mathrm{\mathcal{W}_{1}}\left(\frac{1}{N_{t}}p(\mathcal{X}_{t}),\frac{1}{\sum_{\hat{\mathbf{x}}\in\hat{\mathcal{X}_{t}}}\hat{w}(\hat{\mathbf{x}})}p_{\hat{w}}(\hat{\mathcal{X}_{t}})\right),$ (13)
@@ 4 Experiments
We evaluate VGFM on synthetic and real datasets, and compare it with state-of-the-art approaches, including methods that account for distributional unbalancedness [25, 49, 43] as well as those that assume mass conservation [21, 36, 39]. Our code will be released online.
@@
Synthetic datasets. Inspired by [22, 25], we adopt the Simulation Gene dataset that applies a three-gene regulatory network to produce a quiescent region and an area exhibiting both transition and growth that can be observed. We also use Dyngen [52] to simulate a scRNA-seq dataset from a dynamic cellular process, which exhibits pronounced branching unbalancedness, with significantly different cell abundances across divergent lineages. Following the experiment setup of [30], we use PHATE [53] to reduce its dimensions to 5. Additionally, inspired by [25] that employs a high-dimensional Gaussian mixture model [54] to evaluate the scalability of models, we adopt a more challenging setting: 1000-dimensional Gaussian mixtures.
@@
Real-world dataset. We conduct experiments on two real-world datasets, Embryoid Body (EB) [55] and CITE-seq (CITE) [56], both preprocessed following the procedures in [17, 21]. Each dataset is evaluated under both 5- and 50-dimensional PCA projections. Specifically, the EB (5D), CITE (5D), and CITE (50D) configurations are assessed using a hold-out strategy, same as [36], in which an intermediate time point is excluded during training. The model is then used to predict the distribution at the hold-out time, and we compute the $\mathcal{W}_{1}$ distance between the predicted and true distributions at that time point. The EB (50D) setting is further used for comparison against [25], evaluated using both $\mathcal{W}_{1}$ and the Relative Mass Error (RME) (see below).
Evaluation metrics. We follow [25] to use $\mathcal{W}_{1}$ to measure the distance between the predicted cell distribution and the true cell distribution. Additionaly, to assess the accuracy of the growth function $g_{\omega}$, we normalize the total mass at time 0 to 1, and denote the relative mass at time $t$ w.r.t time 0 as $m_{t}$, i.e., $m_{t}=\frac{N_{t}}{N_{0}}$. The predicted relative mass at time $t$ based on the evolution induced by $g_{\omega}$ is denoted as $\hat{m}_{t}=\sum_{\hat{\mathbf{x}}\in\hat{\mathcal{X}_{t}}}\hat{w}(\hat{\mathbf{x}})/N_{0}$, where $\sum_{\hat{\mathbf{x}}\in\hat{\mathcal{X}_{t}}}\hat{w}(\hat{\mathbf{x}})$ is the predicted total mass at time $t$. We define the relative mass error, denoted as RME, i.e., $\mathrm{RME}=\frac{|m_{t}-\hat{m}_{t}|}{m_{t}}$.
@@
Inplementation details. We employ a 3-layer (5-layer for dimensions greater than 50) MLP with 256 hidden units and LeakyReLU activation to parameterize both the velocity field $v_{\theta}$ and the growth function $g_{\omega}$. Optimization is performed using the Adam optimizer with a learning rate of $10^{-4}$. A warm-up stage of 200-800 iterations is applied (only matching loss in Eq. (12) ), after which the distribution fitting loss $\mathcal{L}_{\mathrm{OT}}$ defined in Eq. (13) is applied for an additional 30 training epochs. We will discuss the strategy to set $\epsilon$ and $\tau$ in the Appendix B.1.

### v2

@@ Abstract
Learning the underlying dynamics of single cells from snapshot data has gained increasing attention in scientific and machine learning research. The destructive measurement technique and cell proliferation/death result in unpaired and unbalanced data between snapshots, making the learning of the underlying dynamics challenging. In this paper, we propose joint Velocity-Growth Flow Matching (VGFM), a novel paradigm that jointly learns state transition and mass growth of single-cell populations via flow matching. VGFM builds an ideal single-cell dynamics containing velocity of state and growth of mass, driven by a presented two-period dynamic understanding of the static semi-relaxed optimal transport, a mathematical tool that seeks the coupling between unpaired and unbalanced data. To enable practical usage, we approximate the ideal dynamics using neural networks, forming our joint velocity and growth matching framework. A distribution fitting loss is also employed in VGFM to further improve the fitting performance for snapshot data. Extensive experimental results on both synthetic and real datasets demonstrate that VGFM can capture the underlying biological dynamics accounting for mass and state variations over time, outperforming existing approaches for single-cell dynamics modeling. Our code is available at https://github.com/DongyiWang-66/VGFM.
@@ 1 Introduction
Inferring the latent dynamics of complex systems from sparse and noisy data is a fundamental challenge in science and engineering. In many domains, e.g., stock markets [gontis2010long], climate systems [franzke2015stochastic], and biological processes [yeo2021generative, schiebinger2019optimal, zhang2021optimal, waddington1942canalization, bunne2023learning, hay2021estimating, oeppen2002broken, zhang2025review, palma2025enforcing], continuous trajectories are rarely fully observed by sensors. Instead, cross-sectional snapshot data collected at discrete time points are commonly provided. This challenge is especially important in single-cell RNA sequencing [macosko2015highly, klein2015droplet, jovic2022single, saliba2014single, haque2017practical], where destructive sampling yields unpaired population-level snapshots across time without having the tracked individual cell fates. Furthermore, due to the mass changes during cellular development or response processes, observed data often exhibits mass unbalancedness across time, violating mass conservation. Consequently, reconstructing the time-evolving, unnormalized density function from limited samples has become an important research problem and attracted increasing attention [jovic2022single, saliba2014single, lahnemann2020eleven].
@@
Deep learning-based models for dynamics inference have demonstrated great potential [tong_trajectorynet_2020, yeo2021generative, bunne2022proximal, bunne2023learning, zhang2025modeling, sun2025variational, tong2024improving, sha2024reconstructing, jiang2024physics, terpin2024learning, zhang2024learning, zhang2024trajectory, gu2025partially, rohbeck2025modeling]. These models typically employ ordinary or stochastic differential equations (ODEs or SDEs) parameterized by neural networks to approximate the velocity field governing density evolution. One class of methods, based on simulation [chen_neural_2018, tong_trajectorynet_2020, yeo2021generative, huguet2022manifold, jiang2024physics, koshizuka2023neural, bunne2022proximal, sha2024reconstructing, zhang2024learning], generates synthetic trajectories by feeding initial data through neural networks and solving ODEs or SDEs numerically, and compares simulated results with observations for loss computation. However, these simulation-based methods rely heavily on numerical solvers during training, which significantly increases computational cost. In high-dimensional settings, the enlarged search space further exacerbates instability, hindering scalability and convergence. Another line of research focuses on simulation-free approaches [lipman_flow_2023, liu2023flow, albergo2023building, pooladian2023multisample, tong2024improving, eyringunbalancedness, kapusniak2024metric, atanackovic2024meta, klein2024genot, tong2024simulation, corso2025composing, rohbeck2025modeling, palmamulti], where the velocity field is trained efficiently by constructing conditional probability paths without trajectory simulation, yielding better efficiency and stability in training compared with simulation approaches.
@@
Nevertheless, the simulation-free approaches mentioned above rely solely on the velocity field $v$, ignoring the unbalancedness of observed data, which can result in incorrect reconstruction of underlying dynamics or unsatisfying generation performance [zhang2024learning, eyringunbalancedness, corso2025composing]. Distributional imbalance is a common phenomenon in single-cell dynamics, where cellular proliferation and death occur, thus necessitating the introduction of a growth term $g$ to allow for mass increasing or decreasing. To address this, [neklyudov2023action, neklyudov2024a] proposed to jointly learn velocity $v$ and growth $g$ by minimizing the Wasserstein-Fisher-Rao (WFR) metric [chizat2018unbalanced, liero2018optimal]. However, it mathematically enforces $v=\nabla g$, which lacks clear mechanistic justification in biological systems. Other WFR-inspired approaches [sha2024reconstructing, zhang2024learning] design separate neural networks for $v$ and $g$ and prioritize fitting the distributions. However, these methods require heavy simulation during training, limiting their scalability to high dimensions.
@@ 2 Background and Related Works
Optimal transport, a tool for transforming one distribution into another at minimal cost [kantorovich1942translocation], has gained remarkable prominence in recent years across fields such as image processing [tang2024residual, tang2025degradation, guoptimal, gushchin2023entropic, korotin2023neural, zheng2025towards] and bioinformatics [tong2024improving, tong_trajectorynet_2020, tong2024simulation, zhang2024learning, sun2025variational]. Given two distributions $p_{0}$ and $p_{1}$ on domain $\Omega$ satisfying $\int_{\Omega}p_{0}(\mathbf{x})\,\mathrm{d}x=\int_{\Omega}p_{1}(\mathbf{x})\,\mathrm{d}x$, the optimal transport [kantorovich1942translocation] aims to find a joint distribution of $p_{0}$ and $p_{1}$, named coupling, such that the transport cost is minimized, formulated as the following optimization problem:
@@
where $c$ is the cost function. For a map $T$, we define $T_{\#}p_{0}(\mathbf{x})=\int_{\mathbf{x}^{\prime}:T(\mathbf{x}^{\prime})=\mathbf{x}}p_{0}(\mathbf{x}^{\prime})\mathop{}\!\mathrm{d}\mathbf{x}^{\prime}$. Let ${\mbox{P}}^{i}(\mathbf{x}_{0},\mathbf{x}_{1})$ $\mathbf{x}_{i}$ for $i=0,1$, then we have ${\rm P}_{\#}^{i}\pi(\mathbf{x}_{i})=\int_{\Omega}\pi(\mathbf{x}_{0},\mathbf{x}_{1})\,\mathrm{d}\mathbf{x}_{i}$ for $i=0,1$. When $c(\mathbf{x}_{0},\mathbf{x}_{1})=\|\mathbf{x}_{0}-\mathbf{x}_{1}\|^{2}$, by Brenier’s theorem [brenier1987decomposition], the optimal coupling can be expressed as $\pi^{*}=(\mathrm{Id},T^{*})_{\#}p_{0}$, where $T^{*}$ is called the Monge map.
@@
Flow matching [lipman_flow_2023, liu2023flow, albergo2023building] is a simulation-free approach where the velocity field is trained efficiently by constructing conditional probability paths without trajectory simulation. Building on this foundation, further improvements have been achieved by incorporating optimal transport guidance [pooladian2023multisample, tong2024improving], considering unbalancedness [eyringunbalancedness], accounting for manifold structures [kapusniak2024metric], and approximating Schrödinger bridge via flow and score matching [tong2024simulation], all of which has been applied to model single-cell dynamics [tong2024improving, tong2024simulation, kapusniak2024metric, eyringunbalancedness]. Different from these methods, our method explicitly models the simultaneous matching of both the velocity field and the growth function, driven from our developed ideal joint state transition and mass growth dynamics, which does not require the construction of conditional probability paths like the above approaches.
@@
Data imbalance is prevalent across a variety of domains, such as image synthesis and protein generation, motivating the development of flow matching models that account for unbalanced distributions [eyringunbalancedness, corso2025composing, choi2024scalable]. However, these methods often do not model growth functions, which are essential in single-cell contexts. In the single-cell domain, several approaches have been proposed to model growth dynamics [neklyudov2023action, neklyudov2024a, zhang2021optimal, sha2024reconstructing, pariset2023unbalanced], yet they either fail to leverage the informative variations in cell abundance observable from snapshot data [neklyudov2023action, neklyudov2024a], or rely heavily on computationally expensive simulations [zhang2021optimal, sha2024reconstructing, pariset2023unbalanced], limiting their scalability and efficiency. Differently, our method explicitly learns the growth function from observed snapshot data, yielding better performance for single cells than these methods as shown in experiments.
@@ 3 Method
Given the snapshot population of single cells, this paper aims to build the dynamic trajectory of single cells. Considering the unbalancedness between snapshot populations due to the undergo cell proliferation or death, we aim to build a model that can transform between unbalanced distributions by jointly learning a velocity field for controlling state transition and a growth function for controlling mass variation. Towards this goal, we propose joint Velocity-Growth Flow Matching (VGFM) based on semi-relaxed optimal transport to learn the dynamics of single cells. As illustrated in Fig. 2, we first present a decoupled understanding of state transition and mass growth for unbalanced dynamics based on semi-relaxed optimal transport (Figs. 2 (a), (b)), upon which we build a dynamic process between unbalanced distributions (Fig. 2 (c) ). Finally, we propose the velocity and growth flow matching using the built dynamic process to learn the velocity and growth with samples (Fig. 2 (d)). Next, we first describe the unbalanced dynamics of single-cell data, and then discuss the details of each component of our method.
@@ 3.3 Velocity and Growth Flow Matching
Since the optimal growth function $g^{*}_{t}$ in Proposition 1 is required to evolve $p_{0}$ to ${\rm P}_{\#}^{0}\pi^{*}$, satisfying the marginal constraint $p_{0}\exp\left(\int_{0}^{\lambda}g^{*}_{s}\,ds\right)=\rm P_{\#}^{0}\pi^{*},$ possible choice for $g^{*}_{t}$ is not unique. For simplicity and ease of implementation, we choose a time-independent form: $g^{*}_{t}(\mathbf{x})=\frac{\log{\rm P}_{\#}^{0}\pi^{*}(\mathbf{x})-\log p_{0}(\mathbf{x})}{\lambda}$, which minimizes the $L^{2}$-norm energy functional of $g$ and can be explained by the Malthusian growth model [Malthus1798], i.e., the exponential growth model (see Appendix A.4). By plugging the expression of $v^{*}_{t}$ and consists with from Proposition 1 and $g_{t}^{*}$ into Eq. (6), we obtain
@@
that can generalize to new samples. We first estimate $\pi^{*}$ and $T^{*}$ using samples. Given samples $\{\mathbf{x}_{0}^{i}\}_{i=1}^{n}$ of $p_{0}$ and $\{\mathbf{x}_{0}^{i}\}_{j=1}^{m}$ of $p_{1}$, We seek the optimal transport plan $\pi^{0\to 1}\approx\pi^{*}(\mathbf{x}_{0},\mathbf{x}_{1})$ by solving the entropy-regularized semi-relaxed transport problem using Sinkhorn-based algorithm [cuturi2013sinkhorn, peyre_computational_2020] as
@@
where $\cdot]_{i}$ is the $i$-th element. The last term in Eq. (12) is because $p_{0}(\mathbf{x}_{0}^{i})=1$, for $i=1,\cdots,n$. In experiments, $\mathcal{L}_{\rm VGFM}$ can be implemented using mini-batch samples. Specifically, we sample $(i,j)$ from $\pi^{0\to 1}$ and $t$ from $\mathcal{U}(0,1)$, then calculate the loss in square brackets. To improve robustness, we add a Gaussian noise to $\mathbf{x}_{t}$ in experiments.
@@ 3.4 Training Process
[…]
To achieve better performance, we incorporate the Wasserstein distance between generated samples and observed samples as part of the loss function. Specifically, let $\mathcal{X}_{t}=\{\mathbf{x}^{i}_{t}\}_{i=1}^{N_{t}}$ for $t=0,1,\cdots,T-1$ denote the observed samples at different time points, and denote $p(\mathcal{X}_{t})=\sum_{i=1}^{N_{t}}\delta_{\mathbf{x}^{i}_{t}}$. We define $\phi_{v_{\theta}}:\mathbb{R}^{d}\times\mathcal{T}\to\mathbb{R}^{d|\mathcal{T}|}$ as the trajectory mapping function parameterized by the neural network $v_{\theta}$, which takes a given starting point as the initial condition and outputs particle coordinates at time indices $\mathcal{T}$ according to ODE dynamics, where $\mathcal{T}$ denotes a set of time steps. Given an initial set $\mathcal{X}_{0}$, the model $\phi_{v_{\theta}}$ predicts particle positions at future time points in $\mathcal{T}$. Specifically, the predicted samples are computed by applying the neural ODE to $\mathcal{X}_{0}$ over the time indices $\{1,\cdots,T-1\}$, i.e., $\hat{\mathcal{X}}_{1},\cdots,\hat{\mathcal{X}}_{T-1}=\phi_{v_{\theta}}(\mathcal{X}_{0},\{1,\cdots,T-1\})$. Similarly, we define $\phi_{g_{\omega}}$ as the particle weight mapping function parameterized by the neural network $g_{\omega}$, which takes the initial weight of particle $i$ as input and outputs the corresponding weight values at $\mathcal{T}$ under ODE dynamics. For simplicity of notation, we use $\hat{w}(\hat{\mathbf{x}})\mbox{ for }\hat{\mathbf{x}}\in\hat{\mathcal{X}}_{t}$ to denote the weight generated by $\phi_{g_{\omega}}$ corresponding to the particles in set $\hat{\mathcal{X}}$ which is generated by $\phi_{v_{\theta}}$ . The distribution fitting loss is defined as the Wasserstein distance as
[…]
@@ 4 Experiments
We evaluate VGFM on synthetic and real datasets, and compare it with state-of-the-art approaches, including methods that account for distributional unbalancedness [zhang2024learning, pariset2023unbalanced, sha2024reconstructing] as well as those that assume mass conservation [tong2024improving, kapusniak2024metric, tong2024simulation]. Our code will be released online.
@@
Synthetic datasets. Inspired by [sha2024reconstructing, zhang2024learning], we adopt the Simulation Gene dataset that applies a three-gene regulatory network to produce a quiescent region and an area exhibiting both transition and growth that can be observed. We also use Dyngen [cannoodt2021spearheading] to simulate a scRNA-seq dataset from a dynamic cellular process, which exhibits pronounced branching unbalancedness, with significantly different cell abundances across divergent lineages. Following the experiment setup of [huguet2022manifold], we use PHATE [moon2019visualizing] to reduce its dimensions to 5. Additionally, inspired by [zhang2024learning] that employs a high-dimensional Gaussian mixture model [ruthotto2020machine] to evaluate the scalability of models, we adopt a more challenging setting: 1000-dimensional Gaussian mixtures.
@@
Real-world dataset. We conduct experiments on three real-world datasets, Embryoid Body (EB) [moon2018manifold], CITE-seq (CITE) [lance2022multimodal] and Pancreas [bastidas2019comprehensive], preprocessed following the procedures in [tong_trajectorynet_2020, tong2024improving, eyringunbalancedness]. EB and CITE dataset are evaluated under both 5- and 50-dimensional PCA projections, while Pancreas dataset is evaluated under 2000 highly variable gene space to assess VGFM’s scalability on real-world data. Specifically, the EB (5D), CITE (5D), and CITE (50D) configurations are assessed using a hold-out strategy, same as [kapusniak2024metric], in which an intermediate time point is excluded during training. The model is then used to predict the distribution at the hold-out time, and we compute the $\mathcal{W}_{1}$ distance between the predicted and true distributions at that time point. The EB (50D) setting is further used for comparison against [zhang2024learning] and ablation study, evaluated using both $\mathcal{W}_{1}$ and the Relative Mass Error (RME) (see below). In the Pancreas dataset, in addition to $\mathcal{W}_{1}$ and RME, we also present the mean–variance trends between generated and real gene expressions and analyze the interpretability of the learned growth function $g$.
[…]
@@
Implementation details. We employ a 3-layer (5-layer for dimensions greater than 50) MLP with 256 hidden units and LeakyReLU activation to parameterize both the velocity field $v_{\theta}$ and the growth function $g_{\omega}$. Optimization is performed using the Adam optimizer with a learning rate of $10^{-3}$ at warm-up stage and $10^{-4}$ after applying distribution fitting loss. A warm-up stage of 500 iterations for synthetic datasets and 5000 iterations for real-world dataset is applied (only matching loss in Eq. (12) ), after which the distribution fitting loss $\mathcal{L}_{\mathrm{OT}}$ defined in Eq. (13) is applied for an additional 30 training epochs. We will discuss the strategy to set $\epsilon$ and $\tau$ in the Appendix B.1.